# Worst-Case to Expander-Case Reductions: Derandomized and Generalized

### v1

@@ 6 Direct-WTERs for Max-Cut, Densest Subgraph, and Graphical OMv
In this section, we begin by further developing this line of research by providing Direct-WTERs for additional problems not considered in [3]. Let us begin with the Max-Cut problem.
@@ 6.1 Direct-WTER for Max-Cut
In the Max-Cut problem, the goal is to compute the maximum cut in a graph $G$, which we denote by $MC(G):=\max_{S\subseteq V}e(S,V\setminus S)$. To make the graph an $\Omega(1)$-expander, simply applying the core gadget does not work because it might affect the maximum cut in $G_{exp}$ unpredictably. To this end, we add a gadget that ensures that any maximum cut in $G_{exp}$ will separate $L$ from $R$. In addition, we modify the core gadget so that every vertex in $G$ will have the same number of neighbors in $L$ and $R$. These two gadgets together ensure that the vertices in $G$ do not get a preference to be on either side of the cut due to symmetry. Consequently, the maximum cut in $G_{exp}$ induces a maximum cut in $G$.
Let us now describe the full reduction. Given $0<\varepsilon\leq 1$, we apply the core gadget with a tradeoff between the conductance and the blowup, as described in Appendix 5.3. Specifically, we augment $G$ to an $\Omega(\varepsilon^{2})$-expander, with a blowup of $2N=\varepsilon n$ in the number of vertices. This is achieved (roughly) by picking $N=\varepsilon/2n$, and adding for every $v\in V$, $\varepsilon^{2}\deg_{G}(v)+3$ neighbors in $L$. To make the construction more symmetric (so it does not affect the maximum cut), we will do the same thing with respect to $R$, i.e., add for $v\in V$ the same number of neighbors as in $L$.
Observe that for every vertex in $L\cup R$, the number of neighbors it has in $V$ is at most $d\leq\frac{2\varepsilon^{2}m+4n}{N}=4\varepsilon m/n+O(1)\leq 4\varepsilon n+O(1)$. In addition, $d$ is also the degree of the expander between $L$ and $R$. Now we add two bi-cliques of size $N\times 3d$ as follows. Add two sets of $F_{L}$ and $F_{R}$ containing $3d$ vertices each, and all $L$-to-$F_{L}$ and $R$-to-$F_{R}$ edges. The purpose of this gadget is to ensure that $L$ and $R$ are separated in a maximum cut. Denote the resulting graph by $G_{exp}$ (see Figure 3).
Figure 3: Our Direct-WTER for the Max-Cut problem.
Observe that the blowup in the number of vertices is $2N+6d\leq 25\varepsilon n+O(1)$. In addition, we claim that the additional gadgets we add to the core gadget, namely the $V$-to-$R$ edges and the bi-cliques, do not ruin the graph’s expansion. In particular, $G_{exp}$ is an $\Omega(\varepsilon^{2})$-expander. To see why, observe that the induced graph on $L\cup F_{R}\cup F_{L}\cup R$ is a bipartite, $\Omega(d)$-edge expander, hence the proof of Lemma 10 holds for it as well (by letting $L\cup R$ play the role of $L$). The next claim shows that the maximum cut in $G_{exp}$ encodes the maximum cut in $G$.
@@ 6.2 DD-WTER for Densest Subgraph
In the Densest Subgraph problem, we define the density of non-empty set $S\subseteq V$ to be $\rho(S):=m_{S}/|S|$, where $m_{S}$ is the number of edges in the subgraph induced by $S$. The goal is to compute $\rho(G):=\max_{S\subseteq V}\rho(S)$. In our DD-WTER, we will make use of the following claim.
@@
In addition, we assume that the density in $G$ is sufficiently large, namely, $m>42n$ [^13] For clarity, we do not attempt to optimize this constant, although it can be reduced further. To make the graph an $\Omega(1)$-expander, we apply the core gadget in a way that only introduces vertices of degree smaller than $m_{exp}/n_{exp}$, where $m_{exp}$ and $n_{exp}$ are the number of edges and vertices in $G_{exp}$, respectively. To this end, consider the graph obtained by applying the core gadget with a tradeoff between the conductance and the blowup, as described in Appendix 5.3. Specifically, we pick the parameters such that in $G_{exp}$, every $v\in V$ is connected to $\lceil\varepsilon\deg_{G}(v)\rceil+3$ vertices in $L$, and $N=|L|=|R|=n(1+o(1))$. The conductance of this graph is $\Omega(\varepsilon)$. We pick $\varepsilon$ to be some sufficiently small constant, which will be determined later.
@@
Proof.
We will show that $m_{exp}/n_{exp}>\mu$, where $\mu$ is the maximum degree in $L$ (and therefore also in $L\cup R$). Hence, by Claim 15, it will follow that any maximum density subgraph in $G_{exp}$ does not contain vertices from $L\cup R$, so it must induce a maximum density subgraph in $G$.
@@
To adapt the Direct-WTER to a DD-WTER, we replace the core gadget with the dynamic core gadget. However, note that in the dynamic core gadget, there are some variations in the degrees due to lazy updates and rebalances. Nonetheless, these variations can be compensated for by picking a smaller $\varepsilon$. To see why, note that in the dynamic graph, we always have $\varepsilon m+3n\leq e(V,L)\leq 4\varepsilon m+4n$. Hence, $m_{exp}\geq(1+2\varepsilon)m+6n$ and $\mu\leq(4\varepsilon m+9n+o(n))/n$. This implies that the above inequality holds when $n$ is sufficiently large, $01-22\varepsilon>$ 0, and $21n/(1-22\varepsilon)<m$. By picking $\varepsilon=1/44$ we obtain the result.
@@ 6.3 DD-WTER for Graphical OMv instances
To make such instances $\Omega(1)$-expanders, we apply the bipartiteness-preserving core-gadget on $G_{M}$ before adding the problem-specific gadgets. In some cases, as we will soon demonstrate, such gadgets preserve expansion, or they can be easily adapted to preserve expansion. Therefore, this adaptation proves that their OMv-hard instances are $\Omega(1)$-expanders. Let us demonstrate this technique for the problem of $st$-SP (see Appendix D for its definition).
@@
We will prove the proposition for an easier problem called $st$-SP ($3$ vs. $5$), where the goal is only to distinguish between $dist(s,t)=3$ and $dist(s,t)\geq 5$. Henzinger et al. [24] proved a lower bound to this problem via a reduction from OuMv: construct $G_{M}$, add vertices $s$ and $t$ to $G_{M}$, and then update the edges between $s$-to-$A$ and $t$-to-$B$ according to the input vectors. It then followed that whenever $u^{T}Mv=1$, then $dist(s,t)=3$, and whenever $u^{T}Mv=0$, $dist(s,t)\geq 5$. By picking $k=\sqrt{m}$, the graph has $O(m)$ edges, and the lower bound the follows was $m^{1/2-\varepsilon}$ per update and $m^{1-\varepsilon}$ per query. We now modify their construction as follows.
Given $M$, apply the bipartiteness-preserving core gadget to $G_{M}$ before adding vertices $s$ and $t$. Then, pick a non-edge in the expander $X$, i.e., $xy\notin E(X)$ for some $x\in L$ and $y\in R$ arbitrarily. The purpose of this modification is to ensure that $s$ and $t$ are connected to the graph without introducing a path of length $<5$ between them. Now, we procceed as in [24]; namely, given vectors $u=(u_{1},u_{2},\ldots,u_{k})$ and $v=(v_{1},v_{2},\ldots,v_{k})$, we update the graph by adding the edges $sa_{i}$ iff $u_{i}=1$, and $tb_{j}$ iff $v_{j}=1$. Note that $u^{T}Mv=1$ iff $dist(s,t)=3$, and otherwise $dist(s,t)\geq 5$. In addition, we claim that the graph is an $\Omega(1)$-expander at all times. This follows from the next claim.
@@
If $G$ is a $\phi$-expander for some $\phi>0$, then adding a vertex $v$ to $G$ and connecting it to $\ell\geq 1$ arbitrary vertices, results in a graph $H$ which is a $\phi/4$-expander.
In our setting, we add two vertices to $G_{M}$, and at all times, each of them is connected to at least one vertex in $G_{M}$. Therefore, since $G_{M}$ is a $\phi$-expander for some constant $\phi$, then the resulting graph is a $\phi/16$-expander at all times according to this claim.
Proof of Claim 17.
Consider a non-empty cut $S\subseteq V(H)$. Denote by $S_{V}=S\cap V$, and assume without loss of generality that $vol_{G}(S_{V})\leq vol_{G}(V\setminus S_{V})$ (we can assume so because it holds either for $S$ or its complement). This assumption is useful since we then have $e_{G}(S_{V},V\setminus S_{V})\geq\phi vol_{G}(S_{V})$, as $G$ is a $\phi$-expander. Now, we split into the following cases. If $S$ does not contain $v$, i.e. $S_{V}=S$, then $vol(S)\leq vol_{G}(S)+|S|\leq 2vol_{G}(S)$, therefore $e(S,V(H)\setminus S)\geq e_{G}(S,V\setminus S)\geq\phi/2\cdot vol(S)$. If $S$ contains $v$, then $vol(S)\leq\ell+2vol_{G}(S_{V})$. Now, if $\ell\leq 2|S_{V}|$, then $vol(S)\leq 4vol_{G}(S_{V})$, and as in the previous case we have $e(S,V(H)\setminus S)\geq e_{G}(S_{V},V\setminus S_{V})\geq\phi/4\cdot vol(S)$. If $\ell>2|S_{V}|$, then $e(\{v\},V\setminus S_{V})\geq\ell/2$. Therefore, we get
@@
Hence, assuming the OMv Conjecture, there is no algorithm for $st$-SP ($3$ vs. $5$) on $\Omega(1)$-expanders, whose preprocessing time is polynomial, update time $m^{1/2-\varepsilon}$, and query time $m^{1-\varepsilon}$, for any $\varepsilon>0$. ∎
@@ 7 Derandomized Direct-WTERs for Additional Problems
In the following reduction, we assume w.l.o.g. that the graph contains no isolated vertices, as isolated vertices can be easily dealt with. Begin by taking a copy of $V$, denoted $V_{ind}=\{v_{ind}\mid v\in V\}$, in which all vertices are independent. In addition, add the following set of edges $\{uv_{ind}|uv\in E\}$. Next, apply the core gadget on the graph, but apply it only to $V_{ind}$ instead of $V\cup V_{ind}$. Apply the core gadget on a graph whose vertex set is $V_{ind}$ $V$ and add edges only from $V_{ind}$ to $L$. Denote the resulting graph by $\bar{G}:=(\bar{V},\bar{E})$. See Figure 4. Observe that for every $k\geq 3$, all $k$-cliques in $G$ appear in $\bar{G}$. In addition, any $k$-clique in $\bar{G}$ that is not contained in $G$ must contain $k-1$ vertices in $V$, and a single vertex $w_{ind}\in V_{ind}$. In this case, there is a corresponding $k$-clique in $G$, obtained by replacing $w_{ind}$ with $w$. From this observation, we get the following corollary:
@@
The proof splits into the following cases. Either the cut $S_{V}$, is hit by $Q$, in which case there are many edges to $Q_{ind}$, and the proof continues by combining the proofs of Claim 20 and Lemma 10. Otherwise, by the properties of $Q$, there must be many outer edges inside $V$ that account for $vol(S_{V})$. Then, the proof follows from the previous results.
@@ 9 Acknowledgements
We would like to thank Thatchaphol Saranurak for clarifying a typo in the mentioned lower bound for the Densest Subgraph problem [24, Table 5].
@@ Appendix C Explicit Expander Construction
Let us present an explicit, deterministic construction of bipartite $\Omega(1)$-expanders, which builds upon known constructions from the literature. For this part only, we will need to introduce the notion of spectral expanders, which are more common in the literature about explicit constructions.
@@
In particular, if $\lambda_{2}$ is sufficiently small, say $\lambda_{2}\leq 0.1d$, then $G$ is a $\Omega(d)$-edge expander and by Proposition 8 also an $\Omega(1)$-expander.
@@
According to [10, Lemma 3.1], the eigenvalues of this new bipartite graph correspond to those of $G$ and their additive inverses. Notably, the second largest eigenvalue remains bounded by $\lambda$ through this transformation.
@@ Appendix D Problem Definitions
In this section, we provide the definitions and some background on each problem that appears in the paper.

### v2

@@ 6 Direct-WTERs for Max-Cut, Densest Subgraph, and Graphical OMv
In this section, we develop further this line of research by providing a Direct-WTER for Max-Cut, a DD-WTER for Densest Subgraph, and a DD-WTER for Graphical OMv instances.
@@ 6.1 Direct-WTER for Max-Cut
In the Max-Cut problem, the goal is to compute the maximum cut in a graph $G$, which we denote by $MC(G):=\max_{S\subseteq V}e(S,V\setminus S)$. To make the graph an $\Omega(1)$-expander, simply applying the core gadget does not work because it might affect the maximum cut in $G_{exp}$ unpredictably. To this end, we add a gadget that ensures that any maximum cut in $G_{exp}$ will separate $L$ from $R$. Additionally, we modify the core gadget so that every vertex in $G$ will have the same number of neighbors in $L$ and $R$. These two gadgets together ensure that the vertices in $G$ do not get a preference to be in the part of $L$ or $R$, due to symmetry. Consequently, the maximum cut in $G_{exp}$ induces a maximum cut in $G$.
Let us now describe the reduction in more detail. Given $0<\varepsilon\leq 1$, apply the core gadget with a tradeoff between the conductance and the blowup, as described in Section 5.3, to get an $\Omega(\varepsilon^{2})$-expander with a blowup of $2N=\varepsilon n$ in the number of vertices. This is achieved by picking $N=\varepsilon/2n$ (roughly), and adding for every $v\in V$, $\varepsilon^{2}\deg_{G}(v)+3$ neighbors in $L$. Add a symmetric copy of the $V$-to-$L$ edges between $V$ and $R$ as well.
Observe that for every vertex in $L\cup R$, the number of neighbors it has in $V$ is at most $d\leq\frac{2\varepsilon^{2}m+4n}{N}=4\varepsilon m/n+O(1)\leq 4\varepsilon n+O(1)$, where $d$ is also the degree of the expander between $L$ and $R$. Now, add two bi-cliques of size $N\times 3d$ as follows: add two sets of $F_{L}$ and $F_{R}$ containing $3d$ vertices each, and all $L$-to-$F_{L}$ and $R$-to-$F_{R}$ edges. The purpose of this gadget is to ensure that $L$ and $R$ are separated in a maximum cut. Denote the resulting graph by $G_{exp}$. See Figure Fig. 3.
Figure 3: A Direct-WTER for the Max-Cut problem.
Observe that the blowup in the number of vertices is $2N+6d\leq 25\varepsilon n+O(1)$. Moreover, we claim that the additional gadgets: the $V$-to-$R$ edges and the bi-cliques, do not ruin the graph’s expansion, and that $G_{exp}$ is an $\Omega(\varepsilon^{2})$-expander. To see why, observe that the induced graph on $L\cup F_{R}\cup F_{L}\cup R$ is a bipartite, $\Omega(d)$-edge expander, hence the proof of Lemma 10 holds for it as well (by letting $L\cup R$ play the role of $L$).
Finallly, the next claim shows that the maximum cut in $G_{exp}$ encodes the maximum cut in $G$.
@@ 6.2 DD-WTER for Densest Subgraph
In the Densest Subgraph problem, we define the density of non-empty set $S\subseteq V$ to be $\rho(S):=m_{S}/|S|$, where $m_{S}$ is the number of edges in the subgraph induced by $S$. The goal is to compute $\rho(G):=\max_{S\subseteq V}\rho(S)$. In our DD-WTER, we will make use of the following claim:
@@
In addition, we assume that the density in $G$ is sufficiently large, namely, $m>42n$ [^9] For clarity, we do not attempt to optimize this constant, although it can be reduced. To make the graph an $\Omega(1)$-expander, we apply the core gadget in a way that only introduces vertices of degree smaller than $m_{exp}/n_{exp}$, where $m_{exp}$ and $n_{exp}$ are the number of edges and vertices in $G_{exp}$, respectively. To this end, consider the graph obtained by applying the core gadget with a tradeoff between the conductance and the blowup, as described in Section 5.3. Specifically, we pick the parameters such that in $G_{exp}$, every $v\in V$ is connected to $\lceil\varepsilon\deg_{G}(v)\rceil+3$ vertices in $L$, and $N=|L|=|R|=n(1+o(1))$. The conductance of this graph is $\Omega(\varepsilon)$. The parameter $\varepsilon$ is chosen to be a sufficiently small constant which will be determined later.
@@
Proof of 16.
We will show that $m_{exp}/n_{exp}>\mu$, where $\mu$ is the maximum degree in $L$ (and therefore also in $L\cup R$). Hence, by 15, it will follow that any maximum density subgraph in $G_{exp}$ does not contain vertices from $L\cup R$, so it must induce a maximum density subgraph in $G$.
@@
To adapt the Direct-WTER to a DD-WTER, we replace the core gadget with the dynamic core gadget. However, note that in the dynamic core gadget there are some variations in the degrees due to lazy updates and rebalances. Nonetheless, these variations can be compensated for by picking a smaller $\varepsilon$, specifically $\varepsilon=1/44$ will suffice.
@@ 6.3 DD-WTER for Graphical OMv instances
To make such instances $\Omega(1)$-expanders, we apply the bipartiteness-preserving core-gadget on $G_{M}$ before adding the problem-specific gadgets. In some cases, as we will soon demonstrate for the $st$-SP problem, such gadgets preserve expansion, or they can be easily adapted to preserve expansion. Therefore, this adaptation proves that their OMv-hard instances are $\Omega(1)$-expanders. Let us now demonstrate this technique for the problem of $st$-SP and prove Proposition 5 which states that $st$-SP is OMv-hard on $\Omega(1)$-expanders.
@@
We will prove the proposition for an easier variant of the problem called $st$-SP ($3$ vs. $5$), where the goal is only to distinguish between $dist(s,t)=3$ and $dist(s,t)\geq 5$. Henzinger et al. [23] proved a lower bound to this problem via a reduction from OuMv: construct $G_{M}$, add vertices $s$ and $t$ to $G_{M}$, and then update the edges between $s$-to-$A$ and $t$-to-$B$ according to the input vectors. It then followed that whenever $u^{T}Mv=1$, then $dist(s,t)=3$, and whenever $u^{T}Mv=0$, $dist(s,t)\geq 5$. By picking $k=\sqrt{m}$, the graph has $O(m)$ edges, and the lower bound the follows was $m^{1/2-\varepsilon}$ per update and $m^{1-\varepsilon}$ per query. We now modify their construction as follows.
Given $M$, apply the bipartiteness-preserving core gadget to $G_{M}$ before adding vertices $s$ and $t$. Then, pick a non-edge in the expander $X$, i.e., $xy\notin E(X)$ for some $x\in L$ and $y\in R$ arbitrarily. The purpose of this modification is to ensure that $s$ and $t$ are connected to the graph without introducing a path of length $<5$ between them. Now, we procceed as in [23]; namely, given vectors $u=(u_{1},u_{2},\ldots,u_{k})$ and $v=(v_{1},v_{2},\ldots,v_{k})$, we update the graph by adding the edges $sa_{i}$ iff $u_{i}=1$, and $tb_{j}$ iff $v_{j}=1$. Note that $u^{T}Mv=1$ iff $dist(s,t)=3$, and otherwise $dist(s,t)\geq 5$. In addition, we claim that the graph is an $\Omega(1)$-expander. This follows from the next claim.
@@
If $G$ is a $\phi$-expander for some $\phi>0$, then adding a vertex to $G$ and connecting it arbitrarily to $\ell\geq 1$ vertices, results in an $\phi/4$-expander.
Proof of 17.
Denote the resulting graph by $H$ and the added vertex by $v$. Consider a non-empty cut $S\subseteq V(H)$. Denote by $S_{V}=S\cap V$, and assume without loss of generality that $vol_{G}(S_{V})\leq vol_{G}(V\setminus S_{V})$ (we can assume so because it holds either for $S$ or its complement). This assumption is useful since we then have $e_{G}(S_{V},V\setminus S_{V})\geq\phi vol_{G}(S_{V})$, as $G$ is a $\phi$-expander. Now, we split into the following cases. If $S$ does not contain $v$, i.e. $S_{V}=S$, then $vol(S)\leq vol_{G}(S)+|S|\leq 2vol_{G}(S)$, therefore $e(S,V(H)\setminus S)\geq e_{G}(S,V\setminus S)\geq\phi/2\cdot vol(S)$. If $S$ contains $v$, then $vol(S)\leq\ell+2vol_{G}(S_{V})$. Now, if $\ell\leq 2|S_{V}|$, then $vol(S)\leq 4vol_{G}(S_{V})$, and as in the previous case we have $e(S,V(H)\setminus S)\geq e_{G}(S_{V},V\setminus S_{V})\geq\phi/4\cdot vol(S)$. If $\ell>2|S_{V}|$, then $e(\{v\},V\setminus S_{V})\geq\ell/2$. Therefore, we get
@@
In our setting, we add two vertices to $G_{M}$, and at all times, each of them is connected to at least one vertex in $G_{M}$. Therefore, since $G_{M}$ is a $\phi$-expander for some constant $\phi$, then the resulting graph is a $\phi/16$-expander at all times according to this claim. Hence, assuming the OMv Conjecture, there is no algorithm for $st$-SP ($3$ vs. $5$) on $\Omega(1)$-expanders, whose preprocessing time is polynomial, update time $m^{1/2-\varepsilon}$, and query time $m^{1-\varepsilon}$, for any $\varepsilon>0$. ∎
@@ 7 Derandomized Direct-WTERs for Additional Problems
In the following reduction, we assume w.l.o.g. that the graph contains no isolated vertices, as isolated vertices can be easily dealt with. Begin by taking a copy of $V$, denoted $V_{ind}=\{v_{ind}\mid v\in V\}$, in which all vertices are independent. In addition, add the following set of edges $\{uv_{ind}|uv\in E\}$. Next, apply the core gadget on the graph, but apply it only to $V_{ind}$ instead of $V\cup V_{ind}$. Apply the core gadget on a graph whose vertex set is $V_{ind}$ $V$ and add edges only from $V_{ind}$ to $L$. Denote the resulting graph by $\bar{G}:=(\bar{V},\bar{E})$. See Fig. 4. Observe that for every $k\geq 3$, all $k$-cliques in $G$ appear in $\bar{G}$. In addition, any $k$-clique in $\bar{G}$ that is not contained in $G$ must contain $k-1$ vertices in $V$, and a single vertex $w_{ind}\in V_{ind}$. In this case, there is a corresponding $k$-clique in $G$, obtained by replacing $w_{ind}$ with $w$. From this observation, we get the following corollary:
@@
The proof splits into the following cases. Either the cut $S_{V}$, is hit by $Q$, in which case there are many edges to $Q_{ind}$, and the proof continues by combining the proofs of 20 and Lemma 10. Otherwise, by the properties of $Q$, there must be many outer edges inside $V$ that account for $vol(S_{V})$. Then, the proof follows from the previous results.
@@ 9 Acknowledgements
We would like to thank Thatchaphol Saranurak for clarifying a typo in the mentioned lower bound for the Densest Subgraph problem [23, Table 5].
@@ Appendix C Explicit Expander Construction
In the appendix, we present an explicit, deterministic construction of bipartite $\Omega(1)$-expanders, which builds upon known constructions from the literature. For this part only, we will need to introduce the notion of spectral expanders, which are more common in the literature about explicit constructions.
@@
In particular, if $\lambda_{2}$ is sufficiently small, say $\lambda_{2}\leq 0.1d$, then $G$ is a $\Omega(d)$-edge expander and by 8 also an $\Omega(1)$-expander.
@@
According to [9, Lemma 3.1], the eigenvalues of this new bipartite graph correspond to those of $G$ and their additive inverses. Notably, the second largest eigenvalue remains bounded by $\lambda$ through this transformation.
@@ Appendix D Problem Definitions
In this appendix, we provide the definitions and some background on each problem that appears in the paper.